OpenStudy (anonymous):

PLEASE PLEASE HELP.......!!!! HOW DO YOU FIND THE ITERATED INTEGRAL OF THIS: 0 to 3, 4 to 5 of x(x^2 + 3y)^(1/2) dy dx ?????!!!!!!!!

OpenStudy (dumbcow):

first integrate with respect to y ... treat "x" like a constant substitute u = x^2 +3y du = 3 dy then plug in 4 and 5 for y then integrate with respect to x u = x^2 du = 2x dx here is solution to check your work http://www.wolframalpha.com/input/?i=integrate+x*sqrt%28x^2%2B3y%29+dy+dx+from+y%3D4+to+5%2C+x%3D0+to+3